OpenStudy (unklerhaukus):

What surface area can 2 grains of Platinum cover if the surface is one atom thick

OpenStudy (anonymous):

2 grains Pt →→ 0.01mole If you want to solve this problem, you must know that the radius of Pt is 135 pm (1.35A) The surface area could be [1.35*10^(-10)](^2) * π * 0.01 * 6.02 * 10(^23) square meter (*APPROX.)

OpenStudy (unklerhaukus):

\[2.00[\text{gr}][\text{Pt}]\]\[=2.00[\text{gr}][\text{Pt}]\times\frac{0.0648[\text{g}]}{[\text{gr}]}\times\frac{1[\text{mole}]}{195[\text g][\text {Pt}]}\]\[=6.65\times10^{-4}[\text{mole}]\]\[=6.65\times10^{-4}[\text{mole}]\times6.022\times10^{23}[\text{atom}/\text{mole}]\]\[=4.00\times10^{20}[\text{atom}]\]\[=4.00\times10^{20}[\text{atom}]\times\pi(1.35[\text Å])^2[\text m^2/\text{atom}]\]\[=22.9[\text m^2]\]